OpenStudy (jiteshmeghwal9):

If the system \[2x+3y-5=0,4x+ky-10=0\] has an infinite number of solutions then:

OpenStudy (jiteshmeghwal9):

Ans. k=6

OpenStudy (anonymous):

For Infinite number of solutions, one condition is there: \[\huge \frac{a_1}{a_2} = \frac{b_1}{b_2} = \frac{c_1}{c_2} \] Use this concept..

OpenStudy (anonymous):

Yes you are right jitesh, k = 6 is right answer..

OpenStudy (jiteshmeghwal9):

But i wants the solution:)

OpenStudy (anonymous):

Your question is firstly incomplete.. This Question says if these equations have Infinite Number of Solutions, then find the value of k...

OpenStudy (jiteshmeghwal9):

ya! i don't see it

OpenStudy (anonymous):

So, here we have to find k.. k = 6 is right..

OpenStudy (jiteshmeghwal9):

i wants solution Plz

OpenStudy (anonymous):

What solution?? I am not getting it..

OpenStudy (amistre64):

if they have infinite number of solutions, then they are the same equation ... and differ by a scalar

OpenStudy (anonymous):

How we found k?? You want that??

OpenStudy (jiteshmeghwal9):

the solution of this question

OpenStudy (amistre64):

2x+3y−5=0 4x+ky−10=0 2x+3y−5=0 2(2x+ky/2−5=0)

OpenStudy (jiteshmeghwal9):

i only wants .how do u get it???

OpenStudy (amistre64):

or 2x+3y−5=0 4x+ky−10=0 2(2x+3y−5=0) 4x+ky−10=0 4x+6y−10=0 4x+ky−10=0

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

I have written the formula above.. Use that: a1 = 2, b1 = 3 a2 = 4 and b2 = ?? Use that: \[\large \color{green}{ \frac{2}{4} = \frac{3}{k} \implies 2k = 12 \implies k = 6}\]

OpenStudy (amistre64):

waters method is fine too :)

OpenStudy (jiteshmeghwal9):

k! i gt it:) Thanx a lot @waterineyes & @amistre64

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

@zepp do you want to see more?? Ha ha ha.. Just kidding..

OpenStudy (zepp):

huh?

OpenStudy (zepp):

I personally prefer @amistre64's method ;P

OpenStudy (anonymous):

Ha ha ha..